ElexusQuimby:

A rescue mission is being conducted using a plane to drop supplies and medical personnel. A helicopter is also sent from the same base station to rescue wounded persons. If the plane flies 400 mph, the helicopter flies 200 mph, and they leave at the same time, how far away is the rescue site if the helicopter arrives 1 hour after the plane arrives? If t represents the time that the plane flies, which of the following equations could be used to solve the problem?

Zackthebest:

https://youtu.be/dUiZnkgshAg

ElexusQuimby:

Help please

grezzyy:

I got

1 attachment
ElexusQuimby:

wth

Narad:

Let distance =d time of flight of plane=t time of flight of helicopter=t+1 d=400t for the plane d= 200(t+1) for the helicopter